ICM 1998, Berlin, Aug. 18–27

Abstracts of Short Communications and Posters

Section:

18. Teaching and Popularization of Mathematics

1991 MS Classification: 00 Rodríguez Del Río, Roberto, Dpto. Matemática Aplicada, Fac. CC Quíimicas, Univ. Complutense de Madrid, Spain, Pacheco Esteban, J. P., Univ. Carlos III de Madrid, Spain, Rodríıguez Trueba, M. I., Univ. Complutense de Madrid, Spain Teaching Calculus with the Computer. An Experience in a Science’s Faculty Teaching Calculus or in general Mathematics with a computer is easy in theory, but difficult in practice. Many problems arise: few PC’s time, how to evaluate the work of students, etc. We present in this Poster part of a teaching experience in Calculus, entirely using Computer, with DERIVE and GNUPLOT in the Faculty of Chemical Sciences, at the Universidad Complutense de Madrid.